OpenStudy (jadenwalker76):

The graph below shows a line segment MN: graph of line MN going through ordered pairs negative 2, 3 and 1, negative 3 Which of the following equations best represents the line segment MN? y = −3x + 2 y = −2x + 1 y = −3x − 2 y = −2x − 1

OpenStudy (peachesncream8):

HI do you think you could post the graph mentioned?

OpenStudy (jadenwalker76):

@peachesncream8 yes

OpenStudy (peachesncream8):

If I read the question right are the ordered pairs: ( -2, 3 ) and ( 1, -3 ) ?

OpenStudy (jadenwalker76):

@peachesncream8

OpenStudy (peachesncream8):

you can use the formula (y-y1) = m (x-x1) plug in both of the points (-2,3) and (1, -3)

OpenStudy (peachesncream8):

once you plug in the the coordinates you can isolate the variable m which is the slope to fill in the equation y=mx+b

OpenStudy (peachesncream8):

so, ( -3 -3 ) = m ( 1+ 2)

OpenStudy (peachesncream8):

then, ( -9 ) = 3m .......... m= -3

OpenStudy (peachesncream8):

so that leaves down to two answer choices: Y= -3x +2 OR Y = -3x -2

OpenStudy (jadenwalker76):

@peachesncream8 the - one?

OpenStudy (peachesncream8):

then substitute in one of the coordinates into the equaation once more and now that you have your slope solve for the equation

OpenStudy (jadenwalker76):

@peachesncream8 okay :)

OpenStudy (peachesncream8):

Are by any chance missing any of the answer choices? I tried solving them as well and came up with the following equations: Y= -3x + 0 and Y= - 3x + 9

OpenStudy (peachesncream8):

Good Luck! : )

OpenStudy (jadenwalker76):

@peachesncream8

OpenStudy (peachesncream8):

You can also try finding the slope which will give you the m in the equation y= mx+b

OpenStudy (peachesncream8):

yea finding the slope would've been easier

OpenStudy (jadenwalker76):

@peachesncream8 thanks alot

OpenStudy (peachesncream8):

your welcome